# MARRS: Masked Autoregressive Unit-based Reaction Synthesis

### v1

@@ Abstract
This work aims at a challenging task: human action-reaction synthesis, i.e., generating human reactions based on the action sequence of the other as conditions. Currently, autoregressive modeling approaches have achieved remarkable performance in motion generation tasks, e.g. text-to-motion. However, vector quantization (VQ) accompanying autoregressive generation has inherent disadvantages, including loss of quantization information, low codebook utilization, etc. Moreover, unlike text-to-motion, which focuses solely on the movement of body joints, human action-reaction synthesis also encompasses fine-grained hand movements. In this work, we propose MARRS, a novel framework designed to generate coordinated and fine-grained reaction motions in continuous representations. Initially, we present the Unit-distinguished Motion Variational AutoEncoder (UD-VAE), which segments the entire body into distinct body and hand units, encoding them independently. Subsequently, we propose Action-Conditioned Fusion (ACF), which involves randomly masking a subset of reactive tokens and extracting specific information about the body and hands from the active tokens. Furthermore, we introduce Adaptive Unit Modulation (AUM) to facilitate interaction between body and hand units by using the information from one unit to adaptively modulate the other. Finally, for the diffusion model, we employ a compact MLP as a noise predictor for each distinct body unit and incorporate the diffusion loss to model the probability distribution of each token. Quantitative and qualitative results demonstrate that our method achieves superior performance. The code will be released upon acceptance.
1 Introduction
@@
In the field of generative computer vision, human-human motion generation holds substantial importance for computer animation [19, 16], game development [28, 1], and robotic control [23, 30, 3, 31]. Existing works primarily treat the actors and reactors equally, often overlooking the challenge of reaction generation in ubiquitous human-human interactions. In this work, we focus on generative models for human action-reaction synthesis, i.e., generating human reactions based on the action sequence of another individual as conditions, which can significantly reduce the workload for animators by enabling them to design an acting character and automatically generate meaningful motions for a reacting character.
@@
Although autoregressive modeling approaches have recently achieved remarkable performance in motion generation tasks [22], it is sub-optimal to apply these methods to human action-reaction synthesis. As shown in Sec. 4.4.4, the performance of reconstruction and generation is unsatisfactory. There are two challenges for this: First, vector quantization (VQ) exhibits inherent limitations. Mapping continuous motion data to a constrained set of discrete tokens inevitably results in a loss of information. Although Momask [9] utilizes successive stages of residual quantization to minimize quantization errors incrementally, complex multistage training does not fundamentally eliminate quantization errors. Also, the notorious codebook collapse [39] makes optimizing the code vectors in the existing VQ-VAE not entirely trivial. Second, lacking the representation for modeling motion details. Previous motion generation methods have primarily concentrated on the movement of body joints, while human action-reaction synthesis not only captures the overall body dynamics, but also incorporates the intricate motions of the hands.
@@
Figure 1: Left: Paradigm comparison of different frameworks. (a) and (b) present the structure of the VQ-VAE-based methods and the structure of the Diffusion-based methods, respectively, while (c) shows the framework of our proposed MARRS. Right: result comparison among our method and other methods on eight metrics.
@@
Recently, some works [12] have explored autoregressive image generation without vector quantization, providing new insight into autoregressive generation methods. In this paper, we propose MARRS (Fig. 1), a novel framework that generates synchronized and fine-grained reaction movements in continuous representations. During the training phase, MARRS follows a two-stage paradigm. In the first stage, for the network to learn the concept of the body and hands, we propose Unit-distinguished Motion Variational AutoEncoder (UD-VAE), which splits the whole-body into body and hands units and then encodes them independently by a variational autoencoder. In the second stage, we propose Action-Conditioned Fusion (ACF), which involves randomly masking out a portion of the reactive tokens and then extracting the information of body and hands from the active tokens separately using the transformer. Next, in order to avoid the lack of mutual perception between the two units (body and hands), we propose Adaptive Unit Modulation (AUM), which utilizes information from the body or hands to modulate the other unit adaptively and enables interaction between the body and hands generators to jointly produce whole-body motion. Finally, we utilize a compact MLP to serve as a noise predictor for each distinct unit, modeling individual tokens separately and integrating the diffusion loss [12] to assess the probability of each token. As shown in Fig. 1, we provide a paradigm comparison of different frameworks and a result comparison among different methods.
@@
  • we propose MARRS, a novel framework that generates synchronized and fine-grained reaction movements. To the best of our knowledge, this is the first successful application of autoregressive generation to the field of action-reaction synthesis.
@@
  • We perform extensive experiments on the benchmark human action-reaction synthesis dataset: NTU120-AS. The results in both online and offline settings, demonstrate the effectiveness of our proposed methods.
2 Related Work
2.1 Motion Generation Framework
Autoencoder-Based Motion Generation TEMOS [21] and T2M [8] use a Transformer-based VAE combined with a text encoder to create motion sequences from textual descriptions. AttT2M [40] and TM2D [7] enhance the learning of discrete latent spaces by integrating a spatial-temporal body-part encoder into VQ-VAE [29], thereby improving expressiveness. T2M-GPT [36] introduces a framework based on VQ-VAE and Generative Pretrained Transformer (GPT) for motion generation. MMM [22] proposes a conditional masked motion architecture based on VQ-VAE. To mitigate quantization errors, MoMask [9] introduces a hierarchical quantization scheme.
@@ 2.1 Motion Generation Framework
Diffusion-Based Motion Generation MotionDiffuse [38] and MDM [27] are groundbreaking frameworks for text-driven motion generation using diffusion models. MotionDiffuse offers several advantageous features, including probabilistic mapping, realistic synthesis, and multi-level manipulation. MDM focuses on directly predicting motion and incorporates a geometric loss to enhance the model’s performance. Instead of using a diffusion model to connect raw motion sequences with conditional inputs, MLD [2] employs a latent diffusion model to significantly reduce training and inference costs. ReMoDiffuse [37] introduces an enhancement mechanism based on dataset retrieval to improve the denoising process in diffusion models.
2.2 Human-Human Motion Generation
@@
Synthesizing human-human interactions is essential for applications in gaming and augmented/virtual reality. Earlier approaches often utilized motion graphs [25] and momentum-based inverse kinematics [11] to simulate the movements of human joints. Recently, more sophisticated techniques have emerged. ComMDM [24] refines a pre-trained text-to-motion diffusion model using a limited dataset of two-person motions. RIG [26] transforms the text describing asymmetric interactions into both active and passive forms to ensure consistent textual context for each participant. To facilitate the interaction by positioning the joints of two individuals accurately, InterControl [33] leverages a Large Language Model to create movement plans through carefully crafted prompts. InterGen [13] introduces a diffusion model with shared parameters and several regularization losses. Designing an efficient approach to modeling between two people, TIMotion [32] proposes a temporal and interactive framework. FreeMotion [5] proposes to integrate the single and multi-person motion by the conditional motion distribution. Based on the premise of actor’s motion, ReGenNet [34] proposes a diffusion-based generative model and explicit distance-based interaction loss.
2.3 Autoregressive Modeling without Quantization
@@
MAR [12] initially utilizes autoregressive models within a continuous-valued domain to represent the probability distribution of each token through a diffusion process. MarDini [14] employs MAR for temporal planning, allowing for video generation based on any number of masked frames at various frame positions. MMAR [35] processes continuous-valued image tokens to prevent information loss and separates the diffusion process from the autoregressive model in the context of multi-modal large language models. MARDM [17] presents a human motion diffusion model capable of executing bidirectional masked autoregression.
3 Method
3.1 Problem Formulation
@@
Figure 2: The overall framework of our proposed MARRS. (a) Whole-body motion is divided into two units: body and hands and then each unit is encoded independently by a VAE. (b) shows the process of the masked reaction generation model. First, the reactive token of each unit obtains the interaction information from the active token through Action-Conditioned Fusion. Then different units acquire the coordinated information through Adaptive Unit Modulation. Finally, we employ a compact MLP to serve as a noise predictor and integrate the diffusion loss to model the probability of each token.
@@
In the context of synthesizing human action-reaction sequences, our objective is to produce a reaction that is conditioned on a given action. Formally, given that the reaction sequence $x^{1:N}=\{x^{i}\}_{i=1}^{N}$ and the action sequence $y^{1:N}=\{y^{i}\}_{i=1}^{N}$, our goal is to model the conditional probability distribution ${\rm P}(x^{1:N}|y^{1:N})$, enabling us to sample realistic reactive motions. As with [34], we utilize the SMPL-X [20] human model to represent the sequence of human motions. Thus, the reaction can be represented as $x^{i}=[\theta^{x}_{i},q^{x}_{i},\gamma^{x}_{i}]$, where $\theta^{x}_{i}\in\mathbb{R}^{3K}$, $q^{x}_{i}\in\mathbb{R}^{3}$, $\gamma^{x}_{i}\in\mathbb{R}^{3}$ are the pose parameters, global orientation, and root translation of the person. Here, $K$ is set to 54 representing the number of body joints, including the jaw, eyeballs, and fingers.
3.2 Overview of MARRS
@@
3.3 Unit-distinguished Motion VAE (UD-VAE)
@@
[…]
The structure of the UD-VAE is shown in Fig. 2. The encoder and decoder adopt 1D ResNet-based network, which is the same as T2M-GPT [36].
3.4 Masked Reaction Generation Model
@@
3.4.1 Action-Conditioned Fusion (ACF)
@@
[…]
@@
3.4.2 Adaptive Unit Modulation (AUM)
The whole-body motion can be divided into two units: the body unit and the hands unit. As noted in [41], depending exclusively on these individual unit motion generators may fail to produce coordinated whole-body motions, as they lack awareness of other units. Hence, we propose AUM, which enables interaction between the body and hands generators to jointly produce whole-body motion.
Given that the reactive token embeddings $\hat{X}_{b}^{fusion}$ and $\hat{X}_{h}^{fusion}$ from Sec. 3.4.1, we utilize the body’s information to adaptively modulate and optimize the position of the hands. The modulated hands embeddings $\hat{X}_{h}^{final}$ are acquired as follows:
@@ 3.4.2 Adaptive Unit Modulation (AUM)
3.4.3 Diffusion for Autoregressive Reaction Generation
Autoregressive methods [22, 9] have shown considerable benefits in motion generation. Instead of modeling the jointed distribution of all tokens, autoregressive models formulate the generation task as “next token prediction”:
@@ 3.4.3 Diffusion for Autoregressive Reaction Generation
As noted in MAR [12], training diffusion models by employing MSE loss to carry out autoregressive generation leads to a disastrous FID score because of the failure of capturing chained probability distributions. Motivated by the successful application of MAR in image generation, we incorporate the diffusion loss [12] to estimate the probability of each token within the field of the reaction generation.
@@
Here $\epsilon_{1}\in\mathbb{R}^{d}$ and $\epsilon_{2}\in\mathbb{R}^{d}$ are Gaussian vectors sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. The noise-corrupted data are ${x}_{b}^{t}=\sqrt{\bar{\alpha}_{t}}{x_{b}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{1}$ and ${x}_{h}^{t}=\sqrt{\bar{\alpha}_{t}}{x_{h}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{2}$, where $\bar{\alpha}_{t}$ is a noise schedule [18] indexed by a time step $t$. We model individual tokens by using a small MLP as a noise predictor for each unit separately. The noise estimators $\epsilon_{\theta_{1}}$ and $\epsilon_{\theta_{2}}$ have the same structure but do not share parameters. The notation $\epsilon_{\theta}({x}_{n}^{t}\,|\,t,{z}_{n})$ means that this network takes ${x}_{n}^{t}$ as the input, and is conditional on both ${t}$ and ${z}_{n}$.
@@
cVAE [10] $77.52^{\pm 7.25}$ $0.899^{\pm 0.0002}$ $10.10^{\pm 0.02}$ $19.38^{\pm 0.16}$ $70.10^{\pm 3.42}$ $0.724^{\pm 0.0002}$ $11.14^{\pm 0.04}$ $18.4^{\pm 0.26}$
AGRoL [4] $38.04^{\pm 1.45}$ $0.932^{\pm 0.0001}$ $10.95^{\pm 0.07}$ $21.44^{\pm 0.34}$ $44.94^{\pm 2.46}$ $0.680^{\pm 0.0001}$ $\underline{12.51^{\pm 0.09}}$ $19.73^{\pm 0.17}$
MDM [27] $40.13^{\pm 3.65}$ $0.955^{\pm 0.0001}$ $\mathbf{10.53^{\pm 0.04}}$ $21.15^{\pm 0.26}$ $54.54^{\pm 3.94}$ $0.704^{\pm 0.0003}$ $11.98^{\pm 0.07}$ $19.45^{\pm 0.20}$
MDM-GRU [27] $5.31^{\pm 0.18}$ $\underline{0.993^{\pm 0.0000}}$ $11.03^{\pm 0.06}$ $25.04^{\pm 0.36}$ $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
ReGenNet [34] $\underline{0.90^{\pm 0.01}}$ $\mathbf{1.000^{\pm 0.0000}}$ $10.69^{\pm 0.05}$ $\underline{26.25^{\pm 0.35}}$ $\underline{11.00^{\pm 0.74}}$ $\underline{0.749^{\pm 0.0002}}$ $13.80^{\pm 0.16}$ $\underline{22.90^{\pm 0.14}}$
@@
Table 1: Comparison to state-of-the-arts on the online, unconstrained setting for human action-reaction synthesis on NTU120-AS [34]. $\pm$ indicates 95% confidence interval, $\rightarrow$ means that closer to Real is better. Bold indicates best result and underline indicates second best.
3.4.4 Inference
Given the active motion sequence, we send them to the encoder of Sec. 3.3 to get active tokens. We expect to generate the entire reactive tokens in $T$ iterations in an autoregressive manner and the mask ratio in each iteration $t$ is determined by a decay scheduling function ${cos}(\frac{\pi t}{2T})$. First, we mask out all the reactive tokens, and feed the active and reactive tokens to ACF and AUM. Second, we take out the tokens that need to be unmasked from the predefined permuted sequences, and send them into the diffusion as conditions. Next, the tokens after sampling are regarded as the new reactive tokens, and the other tokens are masked again until the completion of the $T$ iterations. As for the diffusion procedure, we sample ${x}_{n}^{T}$ from a random Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$. This initial sample is then progressively denoised in a step-by-step manner, transforming ${x}_{n}^{T}$ to ${x}_{n}^{0}$ through the iterative process defined by
[…]
where $\sigma_{t}$ represents the noise level at time step $t$, and $\epsilon$ is sampled from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. Ultimately, the decoder in Sec. 3.3 is used to decode all tokens and convert them back into reaction sequences.
4 Experiments
@@
Upper & Lower [6] $0.718^{\pm 0.2524}$ $\mathbf{1.000^{\pm 0.00}}$ $\underline{10.576^{\pm 0.0479}}$ $26.270^{\pm 0.3119}$ $\underline{10.208^{\pm 0.2587}}$ $0.745^{\pm 0.0002}$ $14.079^{\pm 0.1943}$ $\mathbf{23.794^{\pm 0.2376}}$
6-Unit Division [41] $\mathbf{0.214^{\pm 0.0025}}$ $\mathbf{1.000^{\pm 0.00}}$ $10.503^{\pm 0.0378}$ $\underline{26.522^{\pm 0.3148}}$ $14.965^{\pm 0.8164}$ $\mathbf{0.769^{\pm 0.0001}}$ $\mathbf{13.302^{\pm 0.1301}}$ $23.006^{\pm 0.2964}$
@@
Table 2: Ablation studies on the Unit Division. “w.o. Unit Division” represents keeping whole-body as a complete unit. “Upper & Lower” represents dividing whole-body into the upper-and-lower-body proposed by SCA [6]. “6-Unit Division” represents dividing whole-body into six separate units proposed by ParCo [41]. Our proposed approach achieves a more balanced performance overall under train-conditioned and test-conditioned settings.

### v2

@@ Abstract
This work aims at a challenging task: human action-reaction synthesis, i.e., generating human reactions conditioned on the action sequence of another person. Currently, autoregressive modeling approaches with vector quantization (VQ) have achieved remarkable performance in motion generation tasks. However, VQ has inherent disadvantages, including quantization information loss, low codebook utilization, etc. In addition, while dividing the body into separate units can be beneficial, the computational complexity needs to be considered. Also, the importance of mutual perception among units is often neglected. In this work, we propose MARRS, a novel framework designed to generate coordinated and fine-grained reaction motions using continuous representations. Initially, we present the Unit-distinguished Motion Variational AutoEncoder (UD-VAE), which segments the entire body into distinct body and hand units, encoding each independently. Subsequently, we propose Action-Conditioned Fusion (ACF), which involves randomly masking a subset of reactive tokens and extracting specific information about the body and hands from the active tokens. Furthermore, we introduce Adaptive Unit Modulation (AUM) to facilitate interaction between body and hand units by using the information from one unit to adaptively modulate the other. Finally, for the diffusion model, we employ a compact MLP as a noise predictor for each distinct body unit and incorporate the diffusion loss to model the probability distribution of each token. Both quantitative and qualitative results demonstrate that our method achieves superior performance. The code will be released upon acceptance.
Introduction
@@
In the field of generative computer vision, human-human motion generation holds substantial importance for computer animation (Parent 2012; Magnenat-Thalmann et al. 1985), game development (Urbain 2010; Bethke 2003), and robotic control (Saridis 1983; Wang et al. 2023a; Chen et al. 2023b; Wang et al. 2024a). Existing works primarily treat the actors and reactors equally, often neglecting the challenge of reaction generation in ubiquitous human-human interactions. In this work, we focus on generative models for human action-reaction synthesis, i.e., generating human reactions based on the action sequence of another individual as conditions, which can significantly reduce the workload for animators by enabling them to design an acting character and automatically generate meaningful motions for a reacting character.
@@
Although autoregressive modeling approaches have recently achieved remarkable performance in motion generation tasks (Pinyoanuntapong et al. 2024), it is sub-optimal to apply these methods to human action-reaction synthesis. For example, as shown in Tab. 5 (VQ-VAE and UD-VQ-VAE), the reconstruction and generation performance is unsatisfactory. There are two challenges for this: first, vector quantization (VQ) exhibits inherent limitations. Mapping continuous motion data to a constrained set of discrete tokens inevitably results in a loss of information. Although Momask (Guo et al. 2024) utilizes successive stages of residual quantization to minimize quantization errors incrementally, complex multistage training does not fundamentally eliminate quantization errors. Second, the notorious codebook collapse (Zheng and Vedaldi 2023) makes optimizing the code vectors in the existing VQ-VAE not entirely trivial. Without relying on VQ-VAE, some methods (Shi et al. 2024) have tried autoregressive diffusion in motion generation, but their cumbersome training and inference processes do not achieve stunning performance in single-segment motion generation. Moreover, units splitting has been shown to be an effective strategy in motion generation. However, as the number of units increases, the computational effort of the model also increases. And the importance of mutual perception among units is often neglected. Thus, it is necessary to explore a more rational units splitting as well as to design an efficient and effective units perception method.
@@
Figure 1: Left: Paradigm comparison of different frameworks. (a) and (b) present the structures of the VQ-VAE-based and Diffusion-based methods, respectively, while (c) shows the framework of our proposed MARRS. Right: result comparison among our method and other methods on eight metrics.
@@
Recently, some works (Li et al. 2024) have explored autoregressive image generation without vector quantization, providing new insight into autoregressive generation methods. In this paper, we propose MARRS (Fig. 1), a novel framework that generates synchronized and fine-grained reaction movements in continuous representations. During the training phase, MARRS follows a two-stage paradigm. In the first stage, for the network to learn the concept of the body and hands, we propose Unit-distinguished Motion Variational AutoEncoder (UD-VAE), which splits the whole-body into body and hands units and then encodes them independently using a variational autoencoder. In the second stage, we propose Action-Conditioned Fusion (ACF), which involves randomly masking out a portion of the reactive tokens and then extracting the information of body and hands from the active tokens separately using the transformer. Next, in order to avoid the lack of mutual perception between the two units (body and hands), we propose Adaptive Unit Modulation (AUM), which utilizes information from the body or hands to modulate the other unit adaptively and enables interaction between the body and hands generators to jointly produce whole-body motion. Finally, we utilize a compact MLP to serve as a noise predictor for each distinct unit, modeling individual tokens separately and integrating the diffusion loss (Li et al. 2024) to assess the probability of each token. As shown in Fig. 1, we provide a paradigm comparison of different frameworks and a result comparison among different methods.
@@
  • we propose MARRS, a novel framework that generates synchronized and fine-grained reaction movements. To the best of our knowledge, this is the first successful application of masked autoregressive generation to the field of action-reaction synthesis.
@@
  • We conduct extensive experiments on the benchmark human action-reaction synthesis datasets, NTU120-AS and Chi3D-AS. The results in both online and offline settings demonstrate the effectiveness of our proposed methods.
Related Work
Motion Generation Framework
@@
TEMOS (Petrovich, Black, and Varol 2022) and T2M (Guo et al. 2022) use a Transformer-based VAE combined with a text encoder to create motion sequences from textual descriptions. T2M-GPT (Zhang et al. 2023a) introduces a framework based on VQ-VAE and Generative Pretrained Transformer (GPT) for motion generation. MMM (Pinyoanuntapong et al. 2024) proposes a conditional masked motion architecture based on VQ-VAE. To mitigate quantization errors, MoMask (Guo et al. 2024) introduces a hierarchical quantization scheme. MotionDiffuse (Zhang et al. 2024) and MDM (Tevet et al. 2022) are groundbreaking frameworks for text-driven motion generation using diffusion models. Instead of using a diffusion model to connect raw motion sequences with conditional inputs, MLD (Chen et al. 2023a) employs a latent diffusion model to significantly reduce training and inference costs. ReMoDiffuse (Zhang et al. 2023b) introduces an enhancement mechanism based on dataset retrieval to improve the denoising process in diffusion models.
Human-Human Motion Generation
@@
Synthesizing human-human interactions is essential for applications in gaming and augmented/virtual reality. Earlier approaches often utilized motion graphs (Shum, Komura, and Yamazaki 2007) and momentum-based inverse kinematics (Komura, Ho, and Lau 2005) to simulate the movements of human joints. Recently, more sophisticated techniques have emerged. ComMDM (Shafir et al. 2023) refines a pre-trained text-to-motion diffusion model using a limited dataset of two-person motions. RIG (Tanaka and Fujiwara 2023) transforms the text describing asymmetric interactions into both active and passive forms to ensure consistent textual context for each participant. To facilitate the interaction by positioning the joints of two individuals accurately, InterControl (Wang et al. 2023b) leverages a Large Language Model to create movement plans through carefully crafted prompts. InterGen (Liang et al. 2024) introduces a diffusion model with shared parameters and several regularization losses. Designing an efficient approach to modeling between two people, TIMotion (Wang et al. 2024b) proposes a temporal and interactive framework. FreeMotion (Fan et al. 2024) proposes to integrate the single and multi-person motion by the conditional motion distribution. Based on the premise of actor’s motion, ReGenNet (Xu et al. 2024) proposes a diffusion-based generative model and explicit distance-based interaction loss.
Autoregressive Modeling without Quantization
@@
MAR (Li et al. 2024) initially utilizes autoregressive models within a continuous-valued domain to represent the probability distribution of each token through a diffusion process. MarDini (Liu et al. 2024) employs MAR for temporal planning, allowing for video generation based on any number of masked frames at various frame positions. MMAR (Yang et al. 2024) processes continuous-valued image tokens to prevent information loss and separates the diffusion process from the autoregressive model in the context of multi-modal large language models. AMDM (Shi et al. 2024) utilizes the diffusion model to implement frame-by-frame incremental generation. MARDM (Meng et al. 2024) presents a human motion diffusion model capable of executing bidirectional masked autoregression.
Method
Problem Formulation
@@
Figure 2: The overall framework of our proposed MARRS. (a) Whole-body motion is divided into two units: body and hands and then each unit is encoded independently by a VAE. (b) shows the process of the masked reaction generation model.
@@
In the context of synthesizing human action-reaction sequences, our objective is to produce a reaction that is conditioned on a given action. Formally, given that the reaction sequence $x^{1:N}=\{x^{i}\}_{i=1}^{N}$ and the action sequence $y^{1:N}=\{y^{i}\}_{i=1}^{N}$, our goal is to model the conditional probability distribution ${\rm P}(x^{1:N}|y^{1:N})$, enabling us to sample realistic reactive motions. As with (Xu et al. 2024), we utilize the SMPL-X (Pavlakos et al. 2019) human model to represent the sequence of human motions. Thus, the reaction can be represented as $x^{i}=[\theta^{x}_{i},q^{x}_{i},\gamma^{x}_{i}]$, where $\theta^{x}_{i}\in\mathbb{R}^{3K}$, $q^{x}_{i}\in\mathbb{R}^{3}$, $\gamma^{x}_{i}\in\mathbb{R}^{3}$ are the pose parameters, global orientation, and root translation of the person. Here, $K$ is set to 54 representing the number of body joints, including the jaw, eyeballs, and fingers.
Overview of MARRS
@@
Unit-distinguished Motion VAE (UD-VAE)
@@
$\begin{aligned} \mathcal{L}_{VAE}^{k}&={\rm SmoothL1}(\hat{x}^{1:N}_{k},x^{1:N}_{k})+{\rm SmoothL1}(\hat{y}^{1:N}_{k},y^{1:N}_{k}).\end{aligned}$
The structure of the UD-VAE is shown in Fig. 2. The encoder and decoder adopt 1D ResNet-based network, which is the same as T2M-GPT (Zhang et al. 2023a).
Masked Reaction Generation Model
@@
Action-Conditioned Fusion (ACF).
@@
$\begin{aligned} {\rm Attn}({\rm Q,K,V})=softmax(\frac{({\rm QW^{Q}})\cdot({\rm KW^{K}})^{T}}{\sqrt{C}})\cdot({\rm VW^{V}}),\end{aligned}$
@@
Adaptive Unit Modulation (AUM).
The whole-body motion can be divided into two units: the body unit and the hands unit. As noted in (Zou et al. 2024), depending exclusively on these individual unit motion generators may fail to produce coordinated whole-body motions, as they lack awareness of other units. Hence, we propose AUM, which enables interaction between the body and hands generators to jointly produce whole-body motion.
Given that the reactive token embeddings $\hat{X}_{b}^{fusion}$ and $\hat{X}_{h}^{fusion}$ from ACF, we utilize the body’s information to adaptively modulate and optimize the position of the hands. The modulated hands embeddings $\hat{X}_{h}^{final}$ are acquired as follows:
@@
Diffusion for Autoregressive Reaction Generation.
Autoregressive methods (Pinyoanuntapong et al. 2024; Guo et al. 2024) have shown considerable benefits in motion generation. Instead of modeling the jointed distribution of all tokens, autoregressive models formulate the generation task as “next token prediction”:
@@
As noted in MAR (Li et al. 2024), training diffusion models by employing MSE loss to carry out autoregressive generation leads to a disastrous FID score because of the failure of capturing chained probability distributions. Motivated by the successful application of MAR in image generation, we incorporate the diffusion loss (Li et al. 2024) to estimate the probability of each token within the field of the reaction generation.
@@
Here $\epsilon_{1}\in\mathbb{R}^{d}$ and $\epsilon_{2}\in\mathbb{R}^{d}$ are Gaussian vectors sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. The noise-corrupted data are ${x}_{b}^{t}=\sqrt{\bar{\alpha}_{t}}{x_{b}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{1}$ and ${x}_{h}^{t}=\sqrt{\bar{\alpha}_{t}}{x_{h}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{2}$, where $\bar{\alpha}_{t}$ is a noise schedule (Nichol and Dhariwal 2021) indexed by a time step $t$. We model individual tokens by using a small MLP as a noise predictor for each unit separately. The noise estimators $\epsilon_{\theta_{1}}$ and $\epsilon_{\theta_{2}}$ have the same structure but do not share parameters. The notation $\epsilon_{\theta}({x}_{n}^{t}\,|\,t,{z}_{n})$ means that this network takes ${x}_{n}^{t}$ as the input, and is conditional on both t{t} and ${z}_{n}$.
@@
cVAE (Kingma, Welling et al. 2013) $77.52^{\pm 7.25}$ $0.899^{\pm 0.0002}$ $10.10^{\pm 0.02}$ $19.38^{\pm 0.16}$ $70.10^{\pm 3.42}$ $0.724^{\pm 0.0002}$ $11.14^{\pm 0.04}$ $18.4^{\pm 0.26}$
AGRoL (Du et al. 2023) $38.04^{\pm 1.45}$ $0.932^{\pm 0.0001}$ $10.95^{\pm 0.07}$ $21.44^{\pm 0.34}$ $44.94^{\pm 2.46}$ $0.680^{\pm 0.0001}$ $\underline{12.51^{\pm 0.09}}$ $19.73^{\pm 0.17}$
MDM (Tevet et al. 2022) $40.13^{\pm 3.65}$ $0.955^{\pm 0.0001}$ $\mathbf{10.53^{\pm 0.04}}$ $21.15^{\pm 0.26}$ $54.54^{\pm 3.94}$ $0.704^{\pm 0.0003}$ $11.98^{\pm 0.07}$ $19.45^{\pm 0.20}$
MDM-GRU (Tevet et al. 2022) $5.31^{\pm 0.18}$ $\underline{0.993^{\pm 0.0000}}$ $11.03^{\pm 0.06}$ $25.04^{\pm 0.36}$ $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
ReGenNet (Xu et al. 2024) $\underline{0.90^{\pm 0.01}}$ $\mathbf{1.000^{\pm 0.0000}}$ $10.69^{\pm 0.05}$ $\underline{26.25^{\pm 0.35}}$ $\underline{11.00^{\pm 0.74}}$ $\underline{0.749^{\pm 0.0002}}$ $13.80^{\pm 0.16}$ $\underline{22.90^{\pm 0.14}}$
@@
Table 1: Comparison to state-of-the-arts on the online, unconstrained setting for human action-reaction synthesis on NTU120-AS (Xu et al. 2024). $\pm$ indicates 95% confidence interval, $\rightarrow$ means that closer to Real is better. Bold indicates best result and underline indicates second best.
@@
Upper & Lower (Ghosh et al. 2021) $0.718^{\pm 0.2524}$ $\mathbf{1.000^{\pm 0.00}}$ $\underline{10.576^{\pm 0.0479}}$ $26.270^{\pm 0.3119}$ $\underline{10.208^{\pm 0.2587}}$ $0.745^{\pm 0.0002}$ $14.079^{\pm 0.1943}$ $\mathbf{23.794^{\pm 0.2376}}$
6-Unit Division (Zou et al. 2024) $\mathbf{0.214^{\pm 0.0025}}$ $\mathbf{1.000^{\pm 0.00}}$ $10.503^{\pm 0.0378}$ $\underline{26.522^{\pm 0.3148}}$ $14.965^{\pm 0.8164}$ $\mathbf{0.769^{\pm 0.0001}}$ $\mathbf{13.302^{\pm 0.1301}}$ $23.006^{\pm 0.2964}$
@@
Table 2: Ablation studies on the Unit Division. “w.o. Unit Division” represents keeping whole-body as a complete unit. “Upper & Lower” represents dividing whole-body into the upper-and-lower-body proposed by SCA (Ghosh et al. 2021). “6-Unit Division” represents dividing whole-body into six separate units proposed by ParCo (Zou et al. 2024). Our proposed approach achieves a more balanced performance overall under train-conditioned and test-conditioned settings.
Inference.
Given the active motion sequence, we send them to the encoder of UD-VAE to get active tokens. We expect to generate the entire reactive tokens in $T$ iterations in an autoregressive manner and the mask ratio in each iteration $t$ is determined by a decay scheduling function ${cos}(\frac{\pi t}{2T})$. First, we mask out all the reactive tokens, and feed the active and reactive tokens to ACF and AUM. Second, we take out the tokens that need to be unmasked from the predefined permuted sequences, and send them into the diffusion as conditions. Next, the tokens after sampling are regarded as the new reactive tokens, and the other tokens are masked again until the completion of the $T$ iterations. As for the diffusion procedure, we sample ${x}_{n}^{T}$ from a random Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$. This initial sample is then progressively denoised in a step-by-step manner, transforming ${x}_{n}^{T}$ to ${x}_{n}^{0}$ through the iterative process defined by
$\begin{aligned} {x}_{n}^{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left({x}_{n}^{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left({x}_{n}^{t}|t,z_{n}\right)\right)+\sigma_{t}\epsilon,\end{aligned}$
where $\sigma_{t}$ represents the noise level at time step $t$, and $\epsilon$ is sampled from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. Ultimately, the decoder in UD-VAE is used to decode all tokens and convert them back into reaction sequences. The overall inference flowchart is shown in the supplementary materials.
Experiments